# Covariance Analysis of Impulsive Streaking

### v1

@@
[…]
Covariance Analysis of Impulsive Streaking
Jun Wang1,2,3    Zhaoheng Guo2,3,4    Erik Isele1,2,3    Philip H. Bucksbaum1,2,3    Agostino Marinelli1,3    James P. Cryan1,3    Taran Driver1,3 tdriver@stanford.edu 1Stanford PULSE Institute, SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA 2Department of Applied Physics, Stanford University, Stanford, CA 94305, USA 3SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA 4Paul Scherrer Institute, Villigen, Switzerland
(November 4, 2024)
@@ I Introduction
The motion of electrons in molecules and condensed phase systems takes place on the attosecond timescale. Measuring this motion is critical to our understanding of how matter forms and changes, and will help drive technological development in solar energy harvesting [1, 2] and novel microelectronics [3]. It is now possible to generate pulses and trains of pulses of laser light with sub-femtosecond (i.e. attosecond) duration. These technical developments have launched the field of experimental attosecond science [4]. Yet even with access to attosecond pulses, measuring electron motion with attosecond time resolution is a significant technical challenge. Time-resolved measurements require the ability to synchronize an attosecond pulse with another event, such as a second laser pulse, with attosecond timing precision. One method that has proven highly successful involves synchronizing the attosecond pulse with a longer duration infrared laser pulse, and using the oscillating electric field of the infrared laser pulse to map time onto a measurable quantity such as electron momentum [5]. This measurement has been labeled ‘attosecond streaking’, as the action of the field on the electron is analogous to the action of the time-varying voltage in a streak camera [6, 5]. Since the period of oscillation for an infrared laser is on the few-femtosecond timescale and the action of the field on an electron depends on the phase of the oscillation, this class of experiment has emerged as a powerful probe of attosecond electron dynamics [7], including measurements of photoemission delays, Auger-Meitner decay, and characterization of attosecond pulses  [8, 9, 10, 11].
@@
Importantly, streaking measurements rely on a time reference for the ultrafast process, which can be provided either by the precise timing stability (few-attosecond) between the attosecond pulse and the infrared field, or by a single-shot signal produced in concurrence with the process. The timing stability is straightforwardly achieved when the attosecond pulse has been produced by high harmonic generation (HHG), since the HHG driving field is a natural source of a strong infrared field that is well synchronized to the attosecond pulse. Attosecond x-ray free-electron lasers (XFELs) have many advantageous properties: continuous tunability of photon energy, peak powers that can approach one terawatt, and roughly Fourier-limited pulse durations [12, 13]. The inherent timing instability between an infrared laser and the XFEL pulses can span longer than an infrared period, unlike for high harmonic generation (HHG)-based attosecond pulses. It is desirable to exploit these properties of XFELs while achieving time-resolution shorter than the inherent timing instability, which calls for approaches that employ a concurrent single-shot signal as the time reference. Such approaches have been explored recently, demonstrating the possibility to achieve time resolutions that is below the timing jitters between an infrared laser and XFEL pulses [14, 15, 16]. Specifically, the single-shot reference signal is provided by the photoelectrons produced from direct ionization in the presence of the infrared field. With the duration of the XFEL pulse is much shorter than the infrared laser period, the streaking interaction can be treated in the impulsive regime, in which the streaking laser imparts the same momentum shift to all photoelectrons. The vector momentum shift of these photoelectrons can therefore indicate the x-ray/infrared relative timing on a single-shot basis. In Ref. [15], such a time reference has been employed to achieve attosecond timing resolution for measurements of Auger-Meitner decay by characterizing the photoelectrons’ momentum shift on a single-shot basis [15]. The single-shot quantification of the momentum shift places stringent conditions on the single-shot data quality and risks systematic error. Recently, a correlation-based analysis has been used to extract the photoemission delay in x-ray ionization [17] and measured the delay between two attosecond pulses [16], from streaking measurements performed with several hundred femtoseconds timing instability between the streaking laser and attosecond XFEL pulses. In addition to circumventing the requirement for single-shot analysis, this correlation-based analysis bypasses the requirement for the complex modeling of the streaking process in the retrieval of photoemission delay.
@@
Here we provide a formal treatment for the covariance analysis, focusing on the type of streaking measurements with a circularly polarized infrared dressing field that has been labelled as “angular streaking”. In section II, we introduce our mathematical model for the covariance analysis of impulsive angular streaking, in the presence of large jitter in the x-ray/infrared relative timing. In section III, we describe and compare several methods for extracting the time delay between two impulsive signals. These methods can be applied to measuring relative photoemission delay and/or the delay between two attosecond pulses. In section IV we discuss additional considerations for the interpretation and design of attosecond angular streaking measurements. In section V we generalize our analysis to encode an arbitrary time-dependent signal in the covariance with an impuslive signal in angular streaking.
@@ II Model
To describe the x-ray-matter interaction in the presence of an intense laser-field with vector potential $\bm{A}(t)$, the strong-field approximation can be used to write the probability amplitude for observing a photoelectron with momentum $\bm{p}$ [18]:
@@
In Sec. II.1-IV, we demonstrate the covariance analysis in the situation where $Y$ is also impulsive. The angle $\phi$ between the two streaking vectors, $\bm{k}_{Y}$ and $\bm{k}_{X}=\bm{k}$, corresponds to a time delay $\phi/\omega_{L}$ of the feature $Y$ referenced to $X$. If we assume a counter-clockwise rotation of the vector potential as specified in Eqn. (4), $\phi>0$ means that $Y$ is delayed from $X$. Typically the time delay is shorter than $T_{L}$, much shorter than the evolution of the envelope function $A_{0}(t)$, so the two features share the same magnitude of vector potential $|\bm{A}(t_{s})|$. An interesting consequence of the stationary phase approximation in Eqn. (3) is the difference in streaking amplitudes $|\bm{k}_{X}|{=}k$ and $|\bm{k}_{Y}|$ for two photoelectron features dressed by the same streaking field. Combining Eqn. (4), which describes the vector potential of a circularly polarized dressing field, with the momentum shift from the stationary phase approximation $\bm{k}=e(\bm{A}(t_{s})+\tau\left.\frac{d\bm{A}}{dt}\right|_{t_{s}})$, we find that $|\bm{k}_{Y}|=\lambda k$, where the factor $\lambda\equiv\sqrt{(1+(\omega_{L}\tau_{Y})^{2})/(1+(\omega_{L}\tau_{X})^{2})}$ accounts for the difference between $|\bm{k}_{X}|$ and $|\bm{k}_{Y}|$ resulting from the relative photoemission delay $\tau_{Y}-\tau_{X}$. When $|\tau_{Y}-\tau_{X}|\omega_{L}\ll 1$ rad, this factor is $\lambda\simeq 1$, and as $|\tau_{X}-\tau_{Y}|$ increases, the two streaking amplitudes diverge. In this way, the streaking vector of $Y$ is given by $\bm{k}_{Y}=\lambda k(\cos(\kappa+\phi)\,\bm{e}_{x}+\sin(\kappa+\phi)\,\bm{e}_{y})$.
@@ IV.2 Shot Noise
Figure 10: Effects of counting noise, in the absence of machine fluctuations, on the delay retrieval, with amplitude $k/\Delta p=0.5$ and other conditions kept the same as Fig. 5. At each exposure, $N_{s}=7.2\times 10^{5}$ shots are simulated with shot noise and uniformly distributed $\kappa$. Over 10 repetitions, the mean $\langle\Delta\phi\rangle$ and standard deviation $\sigma[\Delta\phi]$ quantify the systematic error (a) and statistical error (b), respectively, for each method: ND-gradient (green dotted), RR-gradient (red solid), diagonal-agnostic RR-gradient (red dashed), and the gradient-free method (yellow dotted). The inset in (a) shows the K-estimate $\mathrm{K}_{XX}$ at $E_{\mathrm{xp}}=18$. When $\kappa$ is sampled from $\mathcal{U}[0,2\pi]$ instead of empirically uniform, the statistical error of the gradient-based methods asymptotes to the noise floor arising from the $\kappa$-sampling over finite $N_{s}$ (black dot-dashed).
@@
The ND method relies on $\overline{X^{0}},\overline{Y^{0}}$, and $\mathrm{K}_{XY}$, as illustrated in Fig. 4, all of which are free from bias induced by shot noise. Thus the systematic error of the ND method is independent of shot noise but remains at the noiseless limit, as shown by the green trace in Fig. 10(a). The results in Fig. 10 are from the same case as in Fig. 5, with $k/\Delta p=0.5$ fixed. On the other hand, the RR method is affected by the positive bias in part of the K-estimators that arises from shot noise, as visualized in the inset of Fig. 10(a). The impact of the bias can be mitigated by assigning zero weights to these positively biased entries of $\mathrm{K}_{XX}$ in the RR procedure. In simulation, we can set $W_{qq^{\prime}}=1-\delta_{qq^{\prime}}$, since it is straightforward to ensure the independence of the Poisson-sampling process between non-intersecting regions. We refer to this variant of the RR method, which is also free from the systematic errors arising from shot noise, as the diagonal-agnostic rank-reduction (daRR) method. This variation improves the accuracy of the delay extraction in low-count scenarios. In high-count scenarios, the daRR method is less accurate than the RR method, due to the small induced bias in the reconstructed gradient. As shown in Fig. 10(a), the point at which the RR method becomes more accurate than the daRR methods is $E_{\mathrm{xp}}/N_{Q}\approx 10$ for the parameters considered. As $E_{\mathrm{xp}}\to\infty$, the systematic error of the RR method diminishes as $\propto 1/E_{\mathrm{xp}}$, approaching the noiseless limit.
@@
The bias in $\mathrm{K}_{XX}$ and $\mathrm{K}_{YY}$ also affects the gradient-free delay retrieval method. Here, shot noise affects the sample correlation matrix $\mathrm{Corr}_{XY}$ via the positive bias in $(\mathrm{K}_{XX})_{qq}$ and $(\mathrm{K}_{YY})_{pp}$. In the case of $k/\Delta p=0.5,p_{c}/\Delta p=6.6$, the systematic error of the gradient-free method exhibits two plateaus at both the low-count and the high-count ends, with a dip in between, as shown in Fig. 10(a). At high counts we approach the noiseless limit, where the error $\Delta\phi=\phi_{\mathrm{fit}}-\phi$ has the same sign as $\phi$, as shown in Fig. 6. In the low-count region, in contrast, the error $\Delta\phi$ exhibits opposite sign to $\phi$. Between the two plateaus, when the sign of $\Delta\phi$ flips, a minimum systematic error is found. However, the exact $E_{\mathrm{xp}}$ for which the minimal systematic error occurs strongly depends on $k/\Delta p$ and $p_{c}/\Delta p$. Therefore, the flux providing the minimal error can only be accurately identified when a thorough characterization of the system parameters is possible.
@@
The accuracy of the delay retrieval, quantified by the systematic error $\overline{\Delta\phi}$, is insensitive to the number of shots $N_{s}$. This is because the accuracy of $\overline{X^{0}},\overline{Y^{0}}$, and K-estimators, depends on $E_{\mathrm{xp}}$ and not $N_{s}$, when only shot noise is considered. In contrast, precision of the delay retrieval, quantified by the statistical error $\sigma[\Delta\phi]$, depends on $N_{s}$. This is because the variance of $\overline{X^{0}},\overline{Y^{0}}$, and K-estimators scales as ${\propto}1/N_{s}$ when the measurement is repeated over a large number of shots $N_{s}$. Shot noise contributes to the variance in a way that is inversely proportional to the total electron counts detected $N_{s}E_{\mathrm{xp}}$. As shown by the colored traces in Fig. 10(b), with shot noise being the only source of noise, the statistical error of the retrieved delay scales as $\propto 1/\sqrt{N_{s}E_{\mathrm{xp}}}$. In addition to shot noise, in experiments, in the variance of $\overline{X^{0}},\overline{Y^{0}}$, and K-estimators, there is also a contribution from the sampling noise of streaking direction $\kappa$, i.e. the non-uniformity of the distribution of a finite number of measured $\kappa$, which is $\propto 1/N_{s}$ for large $N_{s}$. Therefore the statistical error shrinks as $\sqrt{\mathcal{C}_{\kappa}/N_{s}+\mathcal{C}_{E}/(N_{s}E_{\mathrm{xp}})}$ when both $N_{s}$ and $E_{\mathrm{xp}}$ are large, where $\mathcal{C}_{\kappa}$ and $\mathcal{C}_{E}$ are constants independent of $N_{s}$ and $E_{\mathrm{xp}}$. In the case shown in Fig. 10, these constants are $\mathcal{C}_{E}=2.0\,\mathrm{rad}^{2}$ and $\mathcal{C}_{\kappa}=0.045\,\mathrm{rad}^{2}$. The sampling noise of $\kappa$ does not affect the systematic error of the delay retrieval, because it does not affect the accuracy of $\overline{X^{0}},\overline{Y^{0}}$, and the K-estimators.

### v2

@@
Abstract
We present a comprehensive framework of modeling covariance in angular streaking experiments. Within the impulsive streaking regime, the displacement of electron momentum distribution (MD) provides a tight connection between the dressing-free MD and the dressed MD. Such connection establishes universal structures in the composition of streaking covariance that are common across different MDs, regardless of their exact shape. Building on this robust framework, we have developed methods for retrieving temporal information from angular streaking measurements. By providing a detailed understanding of the covariance structure in angular streaking experiments, our work enables more accurate and robust temporal measurements in a wide range of experimental scenarios.
@@ I Introduction
The motion of electrons in molecules and condensed phase systems takes place on the attosecond timescale. It is now possible to generate light pulses and trains of pulses with sub-femtosecond (i.e. attosecond) duration. These technical developments have launched the field of experimental attosecond science [1]. Even with access to attosecond pulses, measuring electron motion with attosecond time resolution is a significant technical challenge. Time-resolved measurements require the ability to synchronize an attosecond pulse with a second event, such as the interaction with a second laser pulse, with sub-femtosecond precision. One method that has proven highly successful involves combining the attosecond pulse with a longer duration infrared laser pulse, and using the oscillating electric field of the infrared laser pulse to map time onto a measurable quantity such as the momentum of an emitted electron [2]. The term ‘attosecond streaking’ has been coined to describe this class of experiments, as the action of the field on the electron is analogous to the action of the time-varying voltage in a streak camera [3, 2]. Since the period of oscillation of an infrared laser pulse is on the few-femtosecond timescale and the action of the field on an electron depends on the phase of the oscillation, attosecond streaking experiments have emerged as a powerful probe of attosecond electron dynamics [4], including measurements of photoemission delays [5], Auger-Meitner decay [6], and characterization of attosecond pulses  [7, 8].
@@
The principal of laser streaking measurements relies on a time reference for the ultrafast process, which can be provided either by the precise timing stability (few-attosecond) between the attosecond pulse and the infrared field, or through a single-shot self-referencing technique. The timing stability can be achieved when the attosecond pulse has been produced by high harmonic generation (HHG), since the HHG emission is naturally synchronized with the infrared driving field. Meanwhile, attosecond x-ray free-electron lasers (XFELs) have many advantageous properties: continuous tunability of photon energy, peak powers that can approach one terawatt, and roughly Fourier-limited pulse durations [9, 10]. The inherent timing jitter between an infrared laser and the XFEL pulses is typically larger than the infrared period, in contrast to HHG-based sources. To make use of the exceptional properties of XFELs in experiments approaching the attosecond regime, approaches that employ a single-shot time reference signal are required. Such approaches have been explored recently, demonstrating the ability to achieve time resolution better than an optical cycle [11, 12, 13, 14, 15, 16]. The single-shot reference signal can be provided by photoelectrons produced from prompt ionization. When the duration of the XFEL pulse is much shorter than the infrared laser period, the streaking interaction can be treated in the impulsive regime, in which the streaking laser imparts the same momentum shift to all photoelectrons. The momentum shift of the prompt electrons therefore provides a single-shot reference for the relative x-ray/laser arrival time. Such a time reference has been employed to achieve attosecond timing resolution for measurements of Auger-Meitner decay by characterizing the photoelectrons’ momentum shift on a single-shot basis [11, 12, 17]. The single-shot quantification of the momentum shift places stringent conditions on the single-shot data and risks systematic error.
Correlation-based approaches have also been employed to overcome the inherent timing jitter in laser/x-ray measurements [18, 14, 15]. Such a technique has recently been used to extract the photoemission delay in x-ray ionization [16] and measure the delay between two attosecond pulses [13]. Here we provide a formal treatment for this covariance analysis, focusing on measurements with a circularly polarized infrared dressing field, or “angular streaking”. In addition to circumventing the requirement for single-shot analysis, this correlation-based analysis bypasses the requirement for the complex modeling of the streaking process in the retrieval of photoemission delay.
@@
In section II, we introduce our mathematical model for the covariance analysis of impulsive angular streaking, in the presence of large jitter in the x-ray/infrared relative timing. In section III, we describe and compare several methods for extracting the time delay between two impulsive processes. These methods can be used to measure the relative photoemission delay between photoionization events produced by the same pulse, or the delay between attosecond pulses. In section IV we discuss additional considerations for the interpretation and design of attosecond angular streaking measurements. In section V we generalize our analysis to the case of an emission process with a complex (non-impulsive) profile in the time domain.
@@ II Model
To describe the x-ray-matter interaction in the presence of an intense laser-field with vector potential $\bm{A}(t)$, we employ the strong-field approximation to write the probability amplitude for observing a photoelectron with momentum $\bm{p}$ [19]:
@@
In Sec. II.1-IV, we demonstrate the covariance analysis in the situation where $Y$ is also impulsive. The angle $\phi$ between the two streaking vectors, $\bm{k}_{Y}$ and $\bm{k}_{X}=\bm{k}$, corresponds to a time delay $\phi/\omega_{L}$ of the feature $Y$ referenced to $X$. If we assume a counter-clockwise rotation of the vector potential as specified in Eqn. (4), $\phi>0$ means that $Y$ is delayed from $X$. Typically the time delay is shorter than $T_{L}$, much shorter than the evolution of the envelope function $A_{0}(t)$, so the two features share the same magnitude of vector potential $|\bm{A}(t_{s})|$. An interesting consequence of the stationary phase approximation in Eqn. (3) is the difference in streaking amplitudes $|\bm{k}_{X}|{=}k$ and $|\bm{k}_{Y}|$ for two photoelectron features dressed by the same streaking field. Combining Eqn. (4), which describes the vector potential of a circularly polarized dressing field, with the momentum shift from the stationary phase approximation $\bm{k}=e(\bm{A}(t_{s})+\tau\left.\frac{d\bm{A}}{dt}\right|_{t_{s}})$, we find that $|\bm{k}_{Y}|=\lambda k$, where the factor $\lambda\equiv\sqrt{(1+(\omega_{L}\tau_{Y})^{2})/(1+(\omega_{L}\tau_{X})^{2})}$ accounts for the difference between $|\bm{k}_{X}|$ and $|\bm{k}_{Y}|$ resulting from the relative photoemission delay $\tau_{Y}-\tau_{X}$ [23]. When $|\tau_{Y}-\tau_{X}|\omega_{L}\ll 1$ rad, this factor is $\lambda\simeq 1$, and as $|\tau_{X}-\tau_{Y}|$ increases, the two streaking amplitudes diverge. In this way, the streaking vector of $Y$ is given by $\bm{k}_{Y}=\lambda k(\cos(\kappa+\phi)\,\bm{e}_{x}+\sin(\kappa+\phi)\,\bm{e}_{y})$.
@@ IV.2 Shot Noise
Figure 10: Effects of shot noise on the delay retrieval, in the absence of machine fluctuations. Shot noise is introduced to the case shown in Fig. 5. For each level of electron flux, $N_{s}=7.2\times 10^{5}$ shots are simulated with shot noise and uniformly distributed $\kappa$, to test the delay retrieval methods at various $\phi$: ND-gradient (green dotted), RR-gradient (red solid), diagonal-agnostic RR-gradient (red dashed), and the gradient-free method (yellow dotted). Over 10 repetitions, the mean $\overline{\Delta\phi}$ and standard deviation $\sigma[\Delta\phi]$ quantify the systematic error and statistical error, respectively. (a) Systematic error at $k/\Delta p{=}0.5$, of the methods The inset in (a) shows the K-estimate $\mathrm{K}_{XX}$ at $E_{\mathrm{fx}}=18$. (b) Statistical error at $k/\Delta p{=}0.5$. When $\kappa$ is sampled from $\mathcal{U}[0,2\pi]$ instead of strictly uniform, the statistical error of the gradient-based methods asymptotes to a noise floor due to finite $N_{s}$ (black dot-dashed). (c) Total error of the methods $(\overline{(\Delta\phi)^{2}})^{1/2}$ over $k/\Delta p$, at a fixed $E_{\mathrm{fx}}=3\times 10^{2}$. For each method, the triangle marks the minimum.
@@
The ND method relies on $\overline{X^{0}},\overline{Y^{0}}$, and $\mathrm{K}_{XY}$, as illustrated in Fig. 4, all of which are free from bias induced by shot noise. Thus the systematic error of the ND method is independent of shot noise but remains at the noiseless limit, as shown by the green trace in Fig. 10(a). On the other hand, the RR method is affected by the positive bias in part of the K-estimators that arises from shot noise, as visualized in the inset of Fig. 10(a). The impact of the bias can be mitigated by assigning zero weights to these positively biased entries of $\mathrm{K}_{XX}$ in the RR procedure. In simulation, we can set $W_{qq^{\prime}}=1-\delta_{qq^{\prime}}$, since it is straightforward to ensure the independence of the Poisson-sampling process between non-intersecting regions. We refer to this variant of the RR method, which is also free from the systematic errors arising from shot noise, as the diagonal-agnostic rank-reduction (daRR) method. This variation improves the accuracy of the delay extraction in low-count scenarios. In high-count scenarios, the daRR method is less accurate than the RR method, due to the small induced bias in the reconstructed gradient. As shown in Fig. 10(a), the point at which the RR method becomes more accurate than the daRR methods is $E_{\mathrm{fx}}/N_{Q}\approx 10$ for the parameters considered. As $E_{\mathrm{fx}}\to\infty$, the systematic error of the RR method diminishes as $\propto 1/E_{\mathrm{fx}}$, approaching the noiseless limit.
@@
The bias in $\mathrm{K}_{XX}$ and $\mathrm{K}_{YY}$ also affects the gradient-free delay retrieval method. Here, shot noise affects the sample correlation matrix $\mathrm{Corr}_{XY}$ via the positive bias in $(\mathrm{K}_{XX})_{qq}$ and $(\mathrm{K}_{YY})_{pp}$. In the case of $k/\Delta p=0.5,p_{c}/\Delta p=6.6$, the systematic error of the gradient-free method exhibits two plateaus at both the low-count and the high-count ends, with a dip in between, as shown in Fig. 10(a). At high counts we approach the noiseless limit, where the error $\Delta\phi=\phi_{\mathrm{fit}}-\phi$ has the same sign as $\phi$, as shown in Fig. 6. In the low-count region, in contrast, the error $\Delta\phi$ exhibits opposite sign to $\phi$. Between the two plateaus, when the sign of $\Delta\phi$ flips, a minimum systematic error is found. However, the exact $E_{\mathrm{fx}}$ for which the minimal systematic error occurs strongly depends on $k/\Delta p$ and $p_{c}/\Delta p$. Therefore, the flux providing the minimal error can only be accurately identified when a thorough characterization of the system parameters is possible.
@@
The accuracy of the delay retrieval, quantified by the systematic error $\overline{\Delta\phi}$, is insensitive to the number of shots $N_{s}$. This is because the accuracy of $\overline{X^{0}},\overline{Y^{0}}$, and K-estimators, depends on the electron flux and not $N_{s}$, when only shot noise is considered. In contrast, precision of the delay retrieval, quantified by the statistical error $\sigma[\Delta\phi]$, depends on $N_{s}$. This is because the variance of $\overline{X^{0}},\overline{Y^{0}}$, and K-estimators scales as ${\propto}1/N_{s}$ when the measurement is repeated over a large number of shots $N_{s}$. Shot noise contributes to the variance in a way that is inversely proportional to the total electron counts detected $N_{s}E_{\mathrm{fx}}$. As shown by the colored traces in Fig. 10(b), with shot noise alone, the statistical error of the retrieved delay scales as $\propto 1/\sqrt{N_{s}E_{\mathrm{fx}}}$. In addition to shot noise, the variance of $\overline{X^{0}},\overline{Y^{0}}$, and K-estimators, also include the sampling noise of the streaking direction $\kappa$, i.e. the non-uniformity of the distribution of a finite number of measured $\kappa$, which is $\propto 1/N_{s}$ for large $N_{s}$. Therefore the statistical error shrinks as $\sqrt{\mathcal{C}_{\kappa}/N_{s}+\mathcal{C}_{E}/(N_{s}E_{\mathrm{fx}})}$ when both $N_{s}$ and $E_{\mathrm{fx}}$ are large, where $\mathcal{C}_{\kappa}$ and $\mathcal{C}_{E}$ are constants independent of $N_{s}$ and $E_{\mathrm{fx}}$. In the case shown in Fig. 10(a& b), $k/\Delta p=0.5$, and these constants are $\mathcal{C}_{E}=2.0\,\mathrm{rad}^{2}$ and $\mathcal{C}_{\kappa}=0.045\,\mathrm{rad}^{2}$. The sampling noise of $\kappa$ does not introduce systematic error in the delay retrieval, as it preserves the accuracy of $\overline{X^{0}},\overline{Y^{0}}$, and the K-estimators.
The total error of the retrieved delay, defined as $\sqrt{\overline{(\Delta\phi)^{2}}}=\sqrt{(\overline{\Delta\phi})^{2}+\sigma[\Delta\phi]^{2}}$ which combines the systematic and statistical errors, also depends on the streaking amplitude $k$. Figure 10(c) shows this total error varies with $k/\Delta p$ at a fixed electron flux $E_{\mathrm{fx}}=3\times 10^{2}$. At large values of $k/\Delta p$, the total error grows with $k$ due to the increasing systematic errors as discussed in section III. However, as $k/\Delta p$ approaches zero, the total error increases due to the presence of noise. This is because the streaking covariance, which scales as $k^{2}$ in the small amplitude limit (Eqn. 8), is overwhelmed by the noise. The dependence on $k/\Delta p$ shown in Fig. 10(c) represents the typical behavior of the total error in delay retrieval, although the exact optimal streaking amplitude also depends on other parameters such as $E_{\mathrm{fx}}$ and $N_{s}$.